A missile is launched on earth with the initial velocity of 100ms-1 at an elevation of 30 degrees.calculate



1 greatest height attained



2 horizontal range



3 time of fright



4 velocity of a missile at time 8 second

1
Expert's answer
2022-05-13T17:03:38-0400

1.

"H_{\max}=\frac{v_0^2\sin^2\theta}{2g}"

"H_{\max}=\frac{100^2\sin^230^\circ}{2*9.8}=128\:\rm m"

2.

"R=\frac{v_0^2\sin2\theta}{g}"

"R=\frac{100^2\sin60^\circ}{9.8}=884\:\rm m"

3.

"t=\frac{2v_0\sin\theta}{g}"

"t=\frac{2*100\sin30^\circ}{9.8}=10.2\:\rm s"

4.

"v=\sqrt{(v_0\cos\theta)^2+(v_0\sin\theta-gt)^2}"

"v=\sqrt{(100\cos30^\circ)^2+(100\sin30^\circ-9.8*8.0)^2}=91\:\rm m/s"